[Hallc_running] [New Logentry] RC daily update

ciprian at jlab.org ciprian at jlab.org
Thu Oct 7 08:45:02 EDT 2021

Logentry Text:
Accelerator Status:
    Smooth beam delivery. Helicity board will be switched to no delay reporting and remain at 120Hz. Hall A likely to be ready to receive beam tomorrow mid-day. Experts are on standby to execute a fast setup for all 4 halls.

General Hall C Status:
    At 1:30 last night the 15K supply temperature shot up to 16.5 (https://logbooks.jlab.org/entry/3919367). This resulted in no beam for ~20min. Experts are looking into it.
    Dave M pointed out that our ice ball over the target has been growing since Sunday (https://logbooks.jlab.org/entry/3919436).
Hall C PION-LT Data Collection/Analysis Status:
    As of 08:30 today we have accumulated: ~4333 pions (43.33 %) (we collected about 5% of the entire dataset for this kinematic over the owl shift last night)

We will have brief RC Daily Meeting today at 10:00 am to discuss any issues and the plan ahead.

Bluejeans remote connection information:
To join the Meeting:
To join via Room System:
Video Conferencing System: bjn.vc -or-
Meeting ID : 861439920
To join via phone :
1) Dial: +1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers )
2) Enter Conference ID : 861439920 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3919471
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211007/50181905/attachment.html>

More information about the Hallc_running mailing list